# A simple Science  template

### v1

@@
A simple Science  template
First Author1∗†, A. Scientist2†, Someone E. Else2    1Department, Institution, City & Postal Code, Country.    2Another Department, Different Institution, City & Postal Code, Country.    Corresponding author. Email: example@mail.com    These authors contributed equally to this work.
See pages - of manuscript.pdf
Supplementary Materials for
@@
Chen Yu1,∗, David Matthews1,∗ Jingxian Wang1,∗, Jing Gu1, Douglas Blackiston2,3, Michael Rubenstein1, Sam Kriegman1,c
1Center for Robotics and Biosystems, Northwestern University, Evanston, IL, USA.
2Dept. of Biology, Tufts University, Medford, MA, USA.
3Wyss Institute for Biologically Inspired Engineering, Harvard University, Boston, MA, USA.
@@
cCorresponding author. Email: sam.kriegman@northwestern.edu
@@
Other Supplementary Materials for this manuscript:
Movies S1 and S2
@@
To find good metamachine configurations with high locomotive ability, we encoded the combinatorial configuration space into a continuous 8D latent space using a variational autoencoder (VAE). We then searched for good configurations within this compressed latent design space using Bayesian Optimization (BO). We used multilayer perceptrons (MLP) for both the VAE encoder and decoder networks. The size of each hidden layer of the VAE encoder was (512, 128, 64, 64, 16) and the size of each hidden layer of the VAE decoder was (64, 64, 128, 512).
@@
Figure S3: The amputation-agnostic policy in different scenarios. Ten previously-unseen amputations, randomly sampled from each training scenario (one limb removed, two hindlimbs removed, all four limbs removed). For testing, amputated modules are partially removed which is distinct from from training where they were fully removed. Across nearly all tested scenarios, the amputation-agnostic policy (blue to green trajectories) substantially outperforms the original policy (cyan to pink trajectories).
Figure S4: The amputation-agnostic policy against out-of-distribution scenarios. The robustness of the amputation-agnostic policy is evaluated with respect to topologically distinct amputation classes composed of previously unseen sets of amputated modules. For each of six sets of module amputations (two diagonal limbs, and three limbs removed), ten amputation locations were randomly sampled (two rows of robots). Across nearly all damage scenarios, the amputation-agnostic policy (blue to green trajectories) substantially outperformed the original policy (cyan to pink trajectories).

### v2

# See pages - of preprint.docx.pdf

@@
See pages - of preprint.docx.pdf
Supplementary Information for
@@
Chen Yu, David Matthews Jingxian Wang, Jing Gu, Douglas Blackiston, Michael Rubenstein, Sam Kriegman
@@
To whom correspondence should be addressed.
Email: sam.kriegman@northwestern.edu
@@
Other supporting materials for this manuscript include the following:
Movies S1 to S3
@@
S1.3 Caption for Movie S3
Mechatronics of a modular leg. Mechatronic components are assembled into a modular leg.
@@
Each module independently sends observation and telemetry messages to the remote computer, and the remote computer uses the most recent message it receives for each module. If the remote computer fails to receive observation updates from one or more modules, then the action policy may take as input observations from modules at slightly different points in time. Note that use of UDP allows module messages to arrive out of order. Timing information was not provided to the modules, and the modules do not directly communicate with each other. While explicit synchronization is not necessary for the agile dynamic behaviors shown herein, the modules are fully equipped with hardware that can provide distributed and low-latency peer-to-peer communication (ESP-NOW, UWB).
@@
To find good metamachine configurations with high locomotive ability, we encoded the combinatorial configuration space into a continuous 8D latent space using a variational autoencoder (VAE). A VAE was used because it transforms the search problem into a continuous one, which simplifies optimization (52), and because we have considerable experience using this particular encoding. We searched for good configurations within this compressed latent design space using Bayesian Optimization (BO). We used multilayer perceptrons (MLP) for both the VAE encoder and decoder networks. The size of each hidden layer of the VAE encoder was (512, 128, 64, 64, 16) and the size of each hidden layer of the VAE decoder was (64, 64, 128, 512).
@@
Figure S3: The amputation-agnostic policy in different scenarios. Ten previously-unseen amputations, randomly sampled from each training scenario (one limb removed, two hindlimbs removed, all four limbs removed). For testing, amputated modules are partially removed which is distinct from from training where they were fully removed. Across nearly all tested scenarios, the amputation-agnostic policy (blue to green trajectories) substantially outperforms the original policy (cyan to pink trajectories). Denoted below each scenario is the locomotion speed (in m/s) generated by the original (dark purple; bottom left) and amputation agnostic (green; bottom right) policies.
Figure S4: The amputation-agnostic policy against out-of-distribution scenarios. The robustness of the amputation-agnostic policy is evaluated with respect to topologically distinct amputation classes composed of previously unseen sets of amputated modules. For each of six sets of module amputations (two diagonal limbs, and three limbs removed), ten amputation locations were randomly sampled (two rows of robots). Across nearly all damage scenarios, the amputation-agnostic policy (blue to green trajectories) substantially outperformed the original policy (cyan to pink trajectories). Denoted below each scenario is the locomotion speed (in m/s) generated by the original (dark purple; bottom left) and amputation agnostic (green; bottom right) policies.